• Carrie D’s. has 6.2 million shares of common stock outstanding, 2.2 million shares of preferred stock outstanding, and 12.00 thousand bonds. If the common shares sell for $15.20 per share, the preferred shares are selling for $28.20 per share, and the bonds are selling for 108.98 percent of par, what would be the weight used for equity in the computation of Carrie D’s WACC?
Answer #1

Total value of common stock=(6.2*15.2)=$94.24 million

Total value of preferred stock=(2.2*28.2)=$62.04 million

Total value of bonds=12000*(1000*108.98%)=$13.0776 million

Total value=(94.24+62.04+13.0776)=$169.3576 million

Hence weight for equity=(94.24/169.3576)

=0.5565(or 55.65% approx).